# ChatDBG: An AI-Powered Debugging Assistant

### v1

@@ V-B C and C++
Table V summarizes the programs from the BugBench [25] and BugsC++ [26] suites used to evaluate ChatDBG’s effectiveness at debugging unmanaged code. Programs used in this evaluation are all real-world applications with concrete known bugs. The four BugBench programs were selected as the only ones we could retrieve, build and reproduce on our system. The BugsC++ suite does not include the original crash-causing inputs. However, it provides links to the original bug report, CVE identifier, and/or exploit-fixing patch, from which we manually retrieve crash reproduction information. We randomly selected and reproduced four bugs from the “memory error” category.
Some of the studied programs do not crash at run time. We employed AddressSanitizer [28] to force a crash at the moment a memory violations occured to trigger those defects. AddressSanitizer is already capable of reporting some information about the crash when it happens. However, this information is often very dense, and typically points at the symptom of the bug, not its root cause. We did not include that information in the initial prompt.
@@
We run our C/C++ experiments on an an x86 Ubuntu 22.04 server. We use Clang and LLDB 17.0.6 to compile and debug, using flags -g -Og -fno-omit-frame-pointer. ChatDBG used OpenAI’s gpt-4-1106-preview model. Each program was run ten times using queries of the form I am debugging cpp-peglib. Provide the root cause of this crash, for PEG, followed by a request to include code in the response. Average time (27 seconds) and cost (\$$0.060.06$ USD) were comparable to Python.
We manually examined each response to determine if ChatDBG successfully provided an actionable code fix for the proximate cause of the crash or for the underlying root cause. We used the critera outlined in Table V. While fixing root causes is the ultimate goal, fixing proximate causes can still be beneficial as fixing crashes enables further debugging steps.
Figure 8: ChatDBG success rate at fixing the proximate or root cause in C/C++ programs. ChatDBG successfully identified and fixed the root cause 36% of the time and the proximate cause an additional 55% of the time.
@@
V-C Threats to Validity
This paper evaluates ChatDBG on two suites of code. The primary suite is a collection of unpublished student labs that may not be entirely representative of code written by, for example, experienced programmers. The second suite consists of real C/C++ applications and bugs drawn from the BugBench and BugsC++ suites. Unlike the Python suite, the C/C++ source code and the bug fixes for these programs are available on Github, which may lead to data leakage affecting the C/C++ study if those repositories were part of the training set for the LLMs we used. While the C/C++ suite consists of real-world applications, most of the errors are memory errors. Other types of errors, such as assertion failures, concurrency errors, or other logical errors, may lead to different results.
@@ V-C Threats to Validity
Our evaluation depends on a manual and subjective evaluation of whether ChatDBG’s explanation of a bug and its proposed fix are satisfactory. We mitigated the risks of subjective evaluation by using precisely-defined criteria decided upon in advance. Python fixes were deemed successful if the resulting code met the correctness requirements outlined in the assignment. C/C++ fixes were deemed successful at fixing the proximate or root cause using the criteria in Table V. Fixes described in prose were permitted, provided that the details of all necessary changes to the code were made explicit.
VI Conclusion and Future Work
@@
We see several avenues for future work. Incorporating an existing fault localization approach into ChatDBG, rather than relying solely on the LLM’s ability to explore the program’s source code and state, could potentially increase its effectiveness and efficiency by allowing the LLM to focus its attention on suspicious and possibly problematic files, functions, or lines of source code. Similarly, incorporating delta debugging [29] could increase the effectiveness of ChatDBG by limiting the amount of input for an LLM and providing failure-inducing events as guidance. Finally, integrating ChatDBG with a time-travel debugger would expand its reach to exploring program state over time, letting it answer queries that cannot be answered given the current program state.
ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG.

### v2

@@ 5.2. C and C++
Table 5 summarizes the programs extracted from the BugBench (Lu et al., 2005) and BugsC++ (An et al., 2023) suites used to evaluate ChatDBG’s effectiveness at debugging unmanaged code. Programs used in this evaluation are all real-world applications with concrete known bugs. The four BugBench programs were selected as the only ones we could retrieve, build and reproduce on our system. The BugsC++ suite does not include the original crash-causing inputs. However, it provides links to the original bug report, CVE identifier, and/or exploit-fixing patch, from which we manually retrieve crash reproduction information. We randomly selected and reproduced four bugs from the “memory error” category.
Some of the studied programs do not crash at run time. We employed AddressSanitizer (Serebryany et al., 2012) to force a crash at the moment a memory violation occurs to trigger those defects. AddressSanitizer is already capable of reporting some information about the crash when it happens. However, this information is often very dense, and typically points at the symptom of the bug, not its root cause. We did not include that information in the initial prompt.
@@
We ran our C/C++ experiments on an an x86 Ubuntu 22.04 server. We used Clang and LLDB 17 to compile and debug, using flags -g -Og -fno-omit-frame-pointer. ChatDBG used OpenAI’s gpt-4-1106-preview model. Each program was run ten times using queries of the form I am debugging cpp-peglib. Provide the root cause of this crash, for PEG, followed by a request to include code in the response. Average time (27 seconds) and cost (\$$0.060.06$ USD) were comparable to Python.
We manually examined each response to determine if ChatDBG successfully provided an actionable code fix for the proximate cause of the crash or for the underlying root cause. We used the criteria outlined in Table 5. While fixing root causes is the ultimate goal, fixing proximate causes can still be beneficial as fixing crashes enables further debugging steps.
Figure 8. ChatDBG success rate at fixing the proximate or root cause in C/C++ programs. ChatDBG successfully identified and fixed the root cause 36% of the time and the proximate cause an additional 55% of the time.
@@
5.3. Threats to Validity
This paper evaluates ChatDBG on two suites of code. The primary suite is a collection of unpublished student labs that may not be entirely representative of code written by, for example, experienced programmers. The second suite consists of real C/C++ applications and bugs drawn from the BugBench and BugsC++ suites. Unlike the Python suite, the C/C++ source code and the bug fixes for these programs are available on GitHub, which may lead to data leakage affecting the C/C++ study if those repositories were part of the training set for the LLMs we used. While the C/C++ suite consists of real-world applications, most of the errors are memory errors. Other types, such as assertion failures, concurrency errors, or other logical errors, may lead to different results.
@@ 5.3. Threats to Validity
Our evaluation depends on a manual evaluation of whether ChatDBG’s explanation of a bug and its proposed fix are satisfactory. We mitigated the risks of subjectivity by using precisely-defined criteria decided upon in advance. Python fixes were deemed successful if the resulting code met the correctness requirements outlined in the assignment. C/C++ fixes were deemed successful at fixing the proximate or root cause using the criteria in Table 5. Fixes described in prose were permitted, provided that the details of all necessary changes to the code were made explicit.
6. Future Work
We see several promising avenues for future work. Incorporating existing fault localization approaches into ChatDBG, rather than relying solely on the LLM’s ability to explore the program’s code and state, could potentially increase its effectiveness and efficiency by allowing the LLM to focus its attention on suspicious files, functions, or lines of source code. Similarly, incorporating delta debugging (Zeller, 1999) could increase the effectiveness of ChatDBG by limiting the amount of input for an LLM and providing failure-inducing events as guidance. Finally, integrating ChatDBG with a time-travel debugger (Project, 2009; O’Callahan et al., 2017) would expand its reach to exploring program state over time, letting it answer queries that cannot be answered given only the current program state. One challenge of integrating these more sophisticated techniques will be ensuring that the LLM can effectively utilize them, which may necessitate fine tuning or additional training on their usage.
7. Conclusion